OpenStudy (anonymous):

I have a question about the slope.... In the slope formula y=mx+b what is the difference between the y and the b and how do I know when to plug it in to the y or b?

Nnesha (nnesha):

b is y-intercept a point where line cross y-axis(there would be only one ) y = just y-coordinate you don't need to plug anything for y

OpenStudy (anonymous):

Example problem: Line A has slope of -5/3 and passes through the point (-2,7). WHat is the x-intercept of line A?

Nnesha (nnesha):

yes y=mx+b so replace m with -5/3 y=-5/3x+b now you have x y value (-2,7) x=-2 y=7 substitute x and y for their solve for b(y-intercept )

OpenStudy (anonymous):

I just don't understand what to plug into the y or b.... like if they ask it passes through the line then what do i do? use it for the y or b?

Nnesha (nnesha):

x-intercept is a point where line cross the x-axis when y =0 after you find y-intercept substitute y for 0

OpenStudy (anonymous):

Like what are some examples of when i have to use it for y or b? what do they have to say for me to know where to replace it?

Nnesha (nnesha):

when they say passes through the points (x,y) then you have to find b

OpenStudy (anonymous):

so replace the b with the number they say it passes through the points?

OpenStudy (anonymous):

or the y

Nnesha (nnesha):

replace with y not b (b is y-intercept )

OpenStudy (anonymous):

hmmm okay

Nnesha (nnesha):

here is an example slope =2 passes through (4,3) \[\huge\rm \color{reD}{y}=mx+b\]\[(4,\color{ReD}{3})\]

OpenStudy (anonymous):

So in what cases do i replace the b (y-intercept) what kind of problems do they have to ask for

Nnesha (nnesha):

then statement would be like this slope is 3 and ***Y-INTERCEPT =5

Nnesha (nnesha):

they would use the *y-intercept ) word

OpenStudy (anonymous):

hmmm okay

OpenStudy (phi):

maybe this helps??: y = 3x + 2 is a formula to "find y" if you are told x. you then put the x and y together as a package (x,y) for example, if x is 1, y will be 5 (do you see how?) and (1,5) is the "package" which is just a way to say the point (1,5) is on the line. does that part make any sense?

OpenStudy (phi):

the next idea is we can (in theory) find all the points on the line using its formula that can be useful and the reason we want a formula rather than a long boring list of (x,y) pairs

OpenStudy (anonymous):

hmmmm i mean thats easy i'm just hving a hard time understand when to substitute the y or b

Nnesha (nnesha):

 i know why you r confused but y-intercept is the specific point where line cross the y-axis which is also one of the y value but they will always use the y-intercept word for b